# Efficient Reasoning via Thought-Training and Thought-Free Inference

### v1

@@ Abstract
Recent advances in large language models (LLMs) have leveraged explicit Chain-of-Thought (CoT) prompting to improve reasoning accuracy. However, most existing methods primarily compress verbose reasoning outputs. These Long-to-Short transformations aim to improve efficiency, but still rely on explicit reasoning during inference. In this work, we introduce 3TF (Thought-Training and Thought-Free inference), a framework for efficient reasoning that takes a Short-to-Long perspective. We first train a hybrid model that can operate in both reasoning and non-reasoning modes, and then further train it on CoT-annotated data to internalize structured reasoning, while enforcing concise, thought-free outputs at inference time using the no-reasoning mode. Unlike compression-based approaches, 3TF improves the reasoning quality of non-reasoning outputs, enabling models to perform rich internal reasoning implicitly while keeping external outputs short. Empirically, 3TF-trained models obtain large improvements on reasoning benchmarks under thought-free inference, demonstrating that high quality reasoning can be learned and executed implicitly without explicit step-by-step generation.
@@ 1 Introduction
In this work, we propose 3TF, a new framework that approaches efficient reasoning from a Short-to-Long perspective. Instead of compressing explicit reasoning traces, 3TF employs an asymmetric training and inference paradigm: during training, the model learns from CoT-annotated data to internalize reasoning patterns; during inference, it operates in a thought-free mode, directly outputting concise final answers. This design enables the model to retain the benefits of CoT supervision—structured internal reasoning—while avoiding explicit, verbose outputs that cause overthinking.
Unlike prior methods that explicitly constrain reasoning length, 3TF focuses on implicit reasoning transfer. By learning from detailed CoT annotations during training, the model internalizes structured reasoning patterns that can be activated silently at inference, transforming reasoning from an explicit linguistic process into a latent cognitive one. This approach allows the model to adaptively allocate reasoning effort: simple tasks can be solved with minimal internal computation, while complex problems trigger richer latent reasoning without producing verbose intermediate steps.
We implement 3TF on both hybrid reasoning models (HRMs) and base models, demonstrating its effectiveness across architectures with different reasoning priors. These models can flexibly operate between explicit and implicit reasoning modes, allowing concise, thought-free inference while preserving the structured reasoning patterns learned from CoT-annotated training data. This unified design enables controlled reasoning efficiency and adaptability to tasks of varying difficulty, and it naturally supports scaling, as larger models can recover more reasoning capability internally without additional supervision.
@@ 2.2 Efficient Reasoning
Figure 1: Overview of the 3TF Asymmetric Reasoning Paradigm. (Top) The templates used for training and inference. (Bottom) The training and inference flowchart. A Base model is first trained on a mix of Think and No-Think data to create a Hybrid Model. This model is then further fine-tuned using the Think data. For No-Think Inference, the model is prompted with the Inference Template, which suppresses the explicit reasoning trace and compels the model to generate only the final answer.
@@ 3.1 Motivation and Intuition
Recent hybrid reasoning models demonstrate that a single model can flexibly alternate between generating explicit thoughts when necessary and answering directly when possible. This adaptability highlights a growing interest in efficient reasoning that does not always require overt verbalization. However, explicit reasoning often remains verbose, unstable, and computationally expensive.
We propose 3TF (Thought-Training and Thought-Free inference), an asymmetric paradigm that decouples reasoning behavior between training and inference. During training, the model learns from reasoning-augmented data containing both explicit thoughts and final answers; during inference, it is explicitly constrained to suppress reasoning and output only the final answer. Hence, 3TF leverages reasoning supervision during training while enforcing thought-free execution at test time. This asymmetry provides a principled balance between interpretability during learning and efficiency during deployment.
@@ 3.3 Information-Theoretic Explanation
From an information-theoretic perspective, explicit and implicit reasoning correspond to two complementary modes of information transmission between the input $x$ and the final answer $a$: expansion through intermediate reasoning traces and compression through direct answering.
Reasoning as Expansion.
Explicit reasoning unfolds the information contained in $x$ into a higher-bandwidth intermediate representation $t$, forming an information chain $x\rightarrow t\rightarrow a$. Since $t$ captures fine-grained dependencies and latent structure, it generally holds that $I(x;t)\gg I(x;a)$. Thus, reasoning acts as an expansion operator: it temporarily increases the representational capacity so that the subsequent generation of $a$ becomes more deterministic and structured.
Training as Alignment.
During training, the model optimizes the joint process in Equation (1) to maximize the mutual information between input, reasoning, and answer. This can be viewed as maximizing
$\max_{\theta}\;I_{\theta}(x;t)+I_{\theta}(t;a\mid x),$ (3)
which encourages $t$ to preserve input information while being maximally predictive of $a$. Reasoning therefore serves as a structured channel through which the model learns to reorganize and align information before final compression.
Inference as Controlled Compression.
At inference, when reasoning emission is suppressed $(t=\varnothing)$, the model effectively compresses the same latent structure directly into $a$. Since training has already aligned reasoning and answering channels, this compression can maintain high fidelity even in the absence of explicit reasoning tokens. Hence, implicit reasoning operates as an information bottleneck that preserves essential semantics while minimizing representational overhead.
Adaptive Behavior.
Finally, because the information-control mechanism is internally learned, the model dynamically adjusts its implicit reasoning depth according to task complexity: Simple queries lead to minimal compression (short answers), while complex problems elicit richer internal representations even without external traces. In this sense, 3TF realizes a self-regulating, task-conditioned compression policy that unifies interpretability, efficiency, and adaptability in neural reasoning.
@@ 4.1 Experimental Setup
We fine-tune the Qwen3 model using the Swift zhao2024swiftascalablelightweightinfrastructure framework for 3 epochs with a batch size of 32. The learning rate is set to 2e-5 with a warmup ratio of 0.05. We enable sequence packing to improve training efficiency. We use lighteval lighteval for evaluation. The number of samplings during evaluation depends on the dataset: 16 samples per question for AIME 2024, 4 samples per question for MATH500 and Olympiad, and 1 sample per question for GSM8K.
@@ 4.2 Comparison with Baselines
Across all datasets, 3TF attains accuracy comparable to the full Thinking mode while producing substantially shorter responses. For instance, on Qwen3-14B, 3TF achieves 94.4% on GSM8K and 97.8% on MATH500, nearly matching the full-CoT performance (94.6% and 97.8%) but using only 19% and 43% of the output tokens on GSM8K and MATH500, respectively. Similar trends hold across smaller model variants, showing that the proposed asymmetric training–inference design effectively preserves reasoning ability without explicit reasoning traces. Moreover, across datasets of varying difficulty, 3TF demonstrates strong adaptability—maintaining concise outputs on GSM8K while exhibiting a lower compression ratio on more challenging benchmarks such as AIME.
@@ 4.3 Scaling Law of Reasoning Recoverability
These results highlight an important scaling dynamic. Larger reasoning-centric LMs contain stronger latent reasoning priors that can be efficiently unlocked through asymmetric 3TF training. In other words, as scale increases, reasoning ability becomes more recoverable without explicit chain-of-thought. The implication is not simply that scaling improves raw accuracy, but that scaling structurally increases reasoning accessibility. As a result, the efficiency advantage of 3TF compounds with scale: larger models reason better, while also requiring far less explicit reasoning at inference time to maintain that level of quality.
@@ 4.4 Avg@k Results
As shown in Figure 3, 3TF exhibits a consistent scaling trend under Avg@k. Across all four benchmarks, performance improves monotonically with model size, and the gap to Base-Think narrows steadily. On GSM8K and MATH500, 3TF almost matches full CoT at 32B, with residual differences below 1–2%. This suggests that large models can recover most of the reasoning benefit even when inference is executed under no-think.
For harder settings (OLYMPIC, AIME24), smaller models show large degradation relative to Base-Think (up to 30–37% at 4B). Yet scale again closes this gap: at 32B, 3TF recovers more than 10% on OLYMPIC and more than 25% on AIME24. This indicates strong scale-driven reasoning recoverability.
Overall, Avg@k reinforces the same pattern as earlier sections: scale makes reasoning increasingly accessible without explicit chain-of-thought at test time. Practically, 3TF becomes more efficient with scale, paying far fewer tokens for nearly the same effective reasoning competence, while retaining room for sparse targeted reasoning only when required.
@@ 4.5 Effect of Adding No-Think Data
We further evaluate the impact of training with No-Think data on Qwen3-8B. Adding No-Think samples consistently shortens outputs across all datasets (e.g., GSM8K average length drops from 241 to 207 tokens, AIME24 from 7961 to 1642). However, this compression is coupled with substantial accuracy loss: GSM8K drops from 93.0% to 90.3%, and AIME24 from 76.6% to 43.3%. MATH500 and OLYMPIC exhibit the same trend. Figure 4 visualizes the effect of different No-Think ratios.
Even extremely small amounts (1% of training) cause noticeable degradation across all benchmarks. Thus, while No-Think data reliably enforces conciseness, excessive or even minimal inclusion harms reasoning fidelity, highlighting the need for careful balancing of reasoning vs non-reasoning samples during training.
@@ 4.6 Train from Scratch
Unlike previous experiments where 3TF was directly fine-tuned on hybrid reasoning models such as Qwen3, here we first train a base model into a hybrid reasoning form using mixed Think and NoThink data, and then apply 3TF training purely on reasoning samples. This setup examines whether asymmetric supervision remains effective without relying on pre-existing hybrid priors.
As shown in Table 2, 3TF consistently achieves high accuracy with concise outputs. On Qwen2.5-7B-Instruct, it reaches 91.0% on GSM8K and 94.6% on MATH500—comparable to the Hybrid (Think) model but using only about one third of its tokens. Similar gains are observed on Qwen2.5-7B-R1-Distill, where 3TF attains 88.6% and 93.6% accuracy on GSM8K and MATH500 while keeping responses compact. The distilled variant generally yields stronger reasoning recoverability; for example, on AIME24, it achieves 60.0% accuracy with less than half the token length of the Instruct version. Overall, 3TF proves effective even when trained from scratch, showing that asymmetric reasoning fine-tuning can independently induce concise yet faithful reasoning behavior.
@@ 4.7 Analysis of Output Structure
To understand how 3TF alters the internal structure of model outputs, we compare the frequency of canonical “reasoning words” across Think, No-Think, and 3TF. The lexical indicators used as reasoning proxies (e.g., “wait”, “however”, “check”) are listed in the Appendix. Table 3 reports both total counts and per-question frequencies of these words over four reasoning benchmarks.
3TF increases the use of reasoning markers relative to No-Think, but still remains far below full Think mode, often more than an order of magnitude smaller. Thus 3TF does not revert into implicit full chain-of-thought. Instead, it preserves short, targeted inference traces while retaining most of the efficiency gains.
We also observe that as benchmark difficulty and distribution shift increase (e.g., Olympic, AIME), all modes (including 3TF) emit more reasoning markers and produce longer outputs. Models appear to adaptively expose more minimal reasoning only when demanded by harder tasks. Taken together, 3TF induces a sparse reasoning regime: not pure No-Think, but substantially shorter than Think, surfacing only critical reasoning artifacts when needed.

### v3

@@ Abstract
Recent advances in large language models (LLMs) have leveraged explicit Chain-of-Thought (CoT) prompting to improve reasoning accuracy. However, most existing methods primarily focus on compressing verbose reasoning outputs. These Long-to-Short transformations aim to improve efficiency, but require a large amount of short CoT data. In this work, we introduce 3TF (Thought-Training and Thought-Free inference), a framework for efficient reasoning that takes a Short-to-Long perspective. We first train a hybrid model that can operate in both reasoning and non-reasoning modes, and then further train it on CoT-annotated data to internalize structured reasoning, while enforcing concise, thought-free outputs at inference time using the no-reasoning mode. Unlike compression-based approaches, 3TF improves the reasoning quality of non-reasoning outputs, enabling models to perform rich internal reasoning implicitly while keeping external outputs short. Empirically, 3TF-trained models obtain large improvements on reasoning benchmarks under thought-free inference, demonstrating that high quality reasoning can be learned and executed implicitly without explicit step-by-step generation.
@@ 1 Introduction
In this work, we propose 3TF, a new framework that approaches efficient reasoning from a Short-to-Long perspective. Instead of compressing explicit reasoning traces, 3TF employs an asymmetric training and inference paradigm. During training, the model learns from CoT-annotated data to internalize reasoning patterns. During inference, it operates in a thought-free mode, directly outputting concise final answers. This design allows the model to retain the benefits of CoT supervision, such as structured internal reasoning, while avoiding explicit and verbose outputs that lead to overthinking.
Previous SFT approaches typically require access to shortened CoT data for model training, such as summarization via LLMs, manual rewriting, or model merging, which can be costly and time-consuming. Unlike prior methods that explicitly constrain reasoning length, 3TF emphasizes implicit reasoning transfer. The training of 3TF is conducted in two stages: in the first stage, we train a hybrid reasoning model (HRM) capable of flexibly switching between explicit and implicit reasoning modes. In the second stage, the model learns exclusively from detailed CoT data to further strengthen its reasoning capability. The resulting model internalizes structured reasoning patterns, which can be silently activated during inference, effectively transforming reasoning from an explicit linguistic process into a latent cognitive one.
@@ 2.2 Efficient Reasoning
Figure 1: Overview of the 3TF Asymmetric Reasoning Paradigm. (Top) Training and inference templates. (Bottom) Flowchart: A Base model is trained on both Think and No-Think data, then fine-tuned on Think data. For No-Think Inference, the model uses the Inference Template to generate only the final answer.
@@ 3.1 Motivation and Intuition
Recent hybrid reasoning models demonstrate the ability to alternate flexibly between explicit reasoning and direct answers, highlighting a growing interest in efficient reasoning that does not always require overt verbalization. However, explicit reasoning often remains verbose, unstable, and computationally expensive. To address this, we propose 3TF (Thought-Training and Thought-Free inference), an asymmetric paradigm that decouples the reasoning process during training and inference. During training, the model learns from reasoning-augmented data, which includes both explicit reasoning steps and final answers, while at inference time, it is explicitly constrained to suppress reasoning and output only the final answer. This approach enables the model to benefit from reasoning supervision during training, while enforcing thought-free execution during deployment.
@@ 3.3 Information-Theoretic Explanation
From an information-theoretic view, explicit and implicit reasoning are complementary. Explicit reasoning expands input $x$ into intermediate traces $t$ ($x\rightarrow t\rightarrow a$), capturing fine-grained dependencies ($I(x;t)\gg I(x;a)$), while training aligns $t$ with $a$ via
$\max_{\theta}I_{\theta}(x;t)+I_{\theta}(t;a\mid x).$
At inference, suppressing $t$ compresses information directly into $a$, leveraging the aligned channels to retain essential semantics. Explicit reasoning expands and structures information; implicit reasoning compresses it efficiently.
@@ 4.1 Experimental Setup
We fine-tune the Qwen3 model using the Swift zhao2024swiftascalablelightweightinfrastructure framework for 3 epochs with a batch size of 32. The learning rate is set to 2e-5 with a warmup ratio of 0.05. We enable sequence packing to improve training efficiency.  We use LightEval lighteval for evaluation. The number of samplings during evaluation depends on the dataset: 16 samples per question for AIME 2024, 4 samples per question for MATH500 and Olympiad, and 1 sample per question for GSM8K. We compute pass@k based on these sampled generations.
@@ 4.2 Comparison with Baselines
Across all datasets, 3TF attains accuracy comparable to the full Thinking mode while producing substantially shorter responses. For instance, on Qwen3-14B, 3TF achieves 94.4% on GSM8K and 97.8% on MATH500, nearly matching the full-CoT performance but using only 19% and 43% of the output tokens on GSM8K and MATH500, respectively. Similar trends hold across smaller model variants, showing that the proposed asymmetric training–inference design effectively preserves reasoning ability without explicit reasoning traces. 3TF further adapts its output length to task complexity, remaining highly concise on easier datasets while allowing moderately longer responses on more challenging ones.
@@
Takeaway #1 Training on reasoning-rich data alone can markedly improve the performance of non-thinking inference modes while keeping outputs far more concise than full reasoning traces.
@@ 4.3 Scaling Law of Reasoning Recoverability
These results highlight a key scaling dynamic: larger reasoning-centric LMs have stronger latent reasoning priors that can be efficiently unlocked via asymmetric 3TF training. As scale increases, reasoning ability becomes more accessible without explicit chain-of-thought.
Takeaway #2 Scaling up the model steadily narrows the gap between 3TF and full Thinking mode, bringing larger models closer to full reasoning performance.
@@ 4.4 Avg@k Results
As shown in Figure 3, 3TF demonstrates a consistent scaling trend under Avg@k. Across all four benchmarks, performance improves monotonically with model size, narrowing the gap to Base-Think. On GSM8K and MATH500, 3TF at 32B nearly matches full CoT, with differences under 1–2%, suggesting that larger models recover most of the reasoning benefits even without explicit chain-of-thought during inference.
In more challenging settings (OLYMPIC, AIME24), smaller models show significant performance degradation compared to Base-Think (up to 30–37% at 4B). However, scaling mitigates this gap: at 32B, 3TF recovers over 10% on OLYMPIC and more than 25% on AIME24, highlighting the strong recoverability of reasoning through scale.
Takeaway #3 For smaller models, 3TF under Avg@k lags behind the Thinking mode, but this gap steadily diminishes as model scale increases.
@@ 4.5 Effect of Adding No-Think Data
We further evaluate the impact of training with No-Think data on Qwen3-8B. Adding No-Think samples consistently shortens outputs across all datasets (e.g., AIME24 from 7961 to 1642). However, this compression is coupled with substantial accuracy loss: GSM8K drops from 93.0% to 90.3%, and AIME24 from 76.6% to 43.3%. MATH500 and OLYMPIC exhibit the same trend. Figure 4 visualizes the effect of different No-Think ratios. Even extremely small amounts (1% of training) cause noticeable degradation across all benchmarks. Thus, while No-Think data reliably enforces conciseness, we hypothesize that the inclusion of non-thought data causes the model to revert to a purely non-reasoning mode, which undermines its ability to perform complex reasoning tasks and leads to performance degradation.
Takeaway #4 Adding No-Think data significantly reduces 3TF performance.
4.6 Ablation Study on the Datasets
Methods GSM8K@1 MATH-500@4
Acc. Tok. Acc. Tok.
Qwen3-8B
NoThink 89.9 245 93.2 1092
NoThink-T1 90.0 300 91.2 741
NoThink-T2 90.2 250 93.4 1038
3TF 93.0 241 96.6 1358
Table 2: Ablation study on Qwen3-8B comparing different training methods: NoThink, NoThink-T1, NoThink-T2, and 3TF.
In this section, we compare 3TF with two training strategies: NoThink-T1, which uses thought data but excludes reasoning part, and NoThink-T2, which is trained from distilled non-thought data (from Qwen3 Non-Think mode). While NoThink-T1 results in slightly shorter outputs, its performance suffers noticeably compared to 3TF. On the other hand, NoThink-T2, trained directly from non-thought data, shows performance comparable to the baseline non-thought model, with no significant improvements. These results suggest that 3TF effectively combines reasoning supervision with thought-free inference, outperforming methods that prioritize direct fine-tuning with non-thought data.
Takeaway #5 Data containing reasoning improves the performance of No-Think modes, rather than the answer content alone.
@@ 4.7 Train from Scratch
Unlike previous experiments where 3TF was fine-tuned directly on hybrid reasoning models like Qwen3, we first train a base model into a hybrid reasoning form using mixed Think and NoThink data. We then apply 3TF training solely on reasoning samples. This setup tests whether asymmetric supervision can be effective without relying on pre-existing hybrid priors.
As shown in Table 3, 3TF consistently achieves high accuracy with concise outputs. On Qwen2.5-7B-Instruct, it reaches 91.0% on GSM8K and 94.6% on MATH500, comparable to the Hybrid (Think) model while using only about a third of its tokens. Similar improvements are observed on Qwen2.5-7B-R1-Distill, where 3TF achieves 88.6% and 93.6% accuracy on GSM8K and MATH500 with compact responses. The distilled variant shows stronger reasoning recoverability; on AIME24, it reaches 60.0% accuracy with less than half the token length of the Instruct version. Overall, 3TF proves effective even when trained from scratch, demonstrating that asymmetric reasoning fine-tuning can independently induce concise yet faithful reasoning behavior.
@@ 4.8 Analysis of Output Structure
To understand how 3TF alters the internal structure of model outputs, we compare the frequency of canonical "reasoning words" across Think, No-Think, and 3TF. These reasoning proxies (e.g., "wait", "however", "check") are listed in the Appendix, with Table 4 reporting both total counts and per-question frequencies across four reasoning benchmarks. 3TF increases the use of reasoning markers relative to No-Think but remains far below Think, often by more than an order of magnitude. This suggests that 3TF does not fully revert to chain-of-thought reasoning, instead maintaining concise, targeted inference traces for efficiency. As benchmark difficulty increases (e.g., Olympic, AIME), all modes, including 3TF, emit more reasoning markers and generate longer outputs, reflecting adaptive reasoning exposure based on task complexity. Thus, 3TF induces a sparse reasoning regime: it is not pure No-Think but is significantly shorter than Think, only surfacing critical reasoning when needed.
Takeaway #6 3TF introduces limited reasoning into non-thinking mode and adapts: the harder the task, the more reasoning patterns are expressed.